Next Section | Table of Contents


Hole 808I ACORK Operations

After finishing ACORK operations in Hole 1173B, we returned to Hole 808I. Despite the short distance between Sites 1173 and 808, we were unable to transit in DP mode while tripping the pipe because of the current. We arrived back at Hole 808I at 0140 hr on 14 June after a 1.5-hr transit. The seafloor positioning beacon was turned on at 0240 hr. Because of our drill pipe loss at the end of Hole 1173B, our operations began with retrieving 75 additional joints of 5-in drill pipe from the riser hold, making them up into stands, and putting them in the drill pipe racker.

During LWD/MWD drilling earlier in the leg, Hole 808I had been drilled with a 9.875-in bit to 5743 mbrf (1057 mbsf). Our next step in preparing this hole for an ACORK was to open the existing 9.875-in hole to 17.5 in. A drilling BHA was assembled with a 9.875-in pilot bit and a 17.5-in hole opener. The aluminum pipe protectors were removed from six joints of 5.5-in drill pipe so these joints could be used as transition joints in the BHA. Once this was assembled, the bit was lowered to 4380 mbrf, and the VIT camera was deployed at 2115 hr on 14 June; the drill line was slipped and cut while the camera was being lowered. At 2315 hr we resumed lowering the drill string until it reached 4669 mbrf, where it was spaced out for reentry.

We began the search for the reentry cone at 0000 hr on 15 June. Hole 808I was reentered at 0210 hr. Once we were in the hole, the drill pipe was lowered to 4812 mbrf (126 mbsf), and the top drive was picked up at 0345 hr. The VIT camera was retrieved; at 0615 hr we started drilling ahead at 4846 mbrf (160 mbsf). Twenty-barrel mud (sepiolite) sweeps were pumped every 40–50 m. The hole opening continued until we reached a total depth of 975.3 mbsf at 2025 hr on 17 June. We circulated a 40-bbl mud sweep and then started pulling out of the hole. The bit cleared the seafloor at 0730 hr on 18 June.

After the drilling tools were secured, we began moving the ship ~70 nmi upcurrent (west-northwest; ~250°) to begin assembling the ACORK. The transit began at 1600 hr on 18 June. During the transit, the drill floor crew prepared to run the 10.75-in ACORK casing. As soon as the transit ended at 0000 hr on 19 June, we began assembling the ACORK.

The first screen joint was moved to the rig floor; a casing collar was welded to the bottom pin connection to act as a guide shoe. The lowermost packer was made up as the second joint in the string, and a single 0.25-in stainless steel tube was run from the top of the screen to the bottom of the packer. The hydraulic umbilical was started at the top of the lowermost packer. As the casing, screens, and uppermost packer were assembled, hydraulic connections were made at the top and bottom of each screen and packer. Three centralizers were added to each joint, and the umbilical was strapped to the casing near each centralizer with 0.625-in metal banding.

During the remainder of 19 June, 68 additional joints of 10.75-in casing and five more screen joints were added to the ACORK assembly. From 0000 to 0500 hr on 20 June, eight more casing joints, one packer, and the landing joint were added to the string. At 0500 hr the casing running tool, ACORK head, and ported sub assembly were moved to the rig floor and attached to the top of the landing joint. The total length of the ACORK assembly was 964.32 m. The completed ACORK assembly was lowered below the rig floor and hung off the moonpool doors by 0700 hr on 20 June.

The next step was to assemble the drilling BHA that would pass through the ACORK assembly, allowing the bit and underreamer to extend below the guide shoe on the bottom of the ACORK. First, we had to remove the crossover subs that came with the mud motor, as they had 7.625-in instead of 6.625-in connections; an ODP crossover sub connected the mud motor to the 8.25-in drill collar above. We then discovered that the crossovers for the underreamer were incorrectly listed in inventory as 4.5-in regular connections and therefore could not be securely made up to those on the underreamer. Because of the previous BHA losses on this leg, no other appropriate crossovers were available, so we welded one of the existing crossover subs directly to the underreamer. The other crossover required the fabrication of a spacer ring to fill a gap between the mating shoulders before it could be welded to the underreamer. A total of 9 hr was required for these modifications.

The remainder of the drilling BHA was assembled without additional problems. The complete drilling BHA contained the following components: 9.875-in bit, bit sub with float valve, three crossover subs, 17-in underreamer, crossover sub, mud motor, crossover sub, nine 8.25-in drill collars, crossover sub, 91 joints of 5.5-in drill pipe, one 5-ft 5.5-in drill pipe pup joint, crossover sub, landing saver sub, one 8.25-in drill collar pup joint, jet sub, MDCB latch sleeve sub, the casing running tool, two 8.25-in drill collars, one tapered drill collar, six joints 5.5-in drill pipe, and crossover sub. The total length of BHA from the running tool to the bit was 975.43 m.

When the drilling BHA was assembled and the casing running tool latched to the ACORK head, the complete assembly was lowered so that the final umbilical connections could be made in the moonpool area. After the ACORK assembly was dipped into the water for purging air from the hydraulic lines, the ACORK assembly and drilling BHA were lowered until the bit was at 1294 mbrf. At 0400 hr the VIT camera was deployed so that we could verify the space out of the underreamer with respect to the bottom of the casing as well as the functioning of the mud motor and underreamer. Once these tests were finished, the VIT camera was recovered at 0545 hr on 21 June. We then lowered the pipe to 3462.15 mbrf, where it was kept during the transit back to the site in DP positioning mode.

During the entire assembly/testing of the ACORK and the drilling BHA, the ship was in a controlled drift back toward Hole 808I. The initial offset from the site was ~70 nmi at 0000 hr on 19 June. We arrived back at Hole 808I at 0330 hr on 22 June.

On 21 June (0620–0645 hr), the Aso Maru arrived to deliver two engineers to assist with the thermistor string deployment later in the leg. A number of small packages, fresh fruit, and vegetables were also delivered. Two small pieces of failed drill pipe were offloaded for forensic analysis. Also, during the DP transit back to the site, 113 aluminum pipe protectors were removed from the drill pipe, as they would not be used.

We resumed lowering the drill pipe once we were back on site and ran the pipe down to 3798 mbrf, where we stopped to deploy the VIT camera at 0430 hr on 22 June. The pipe trip then continued at 0500 hr. Once the drill string reached 4662 mbrf, it was spaced out in preparation for reentering Hole 808I. After only 20 min, Hole 808I was reentered at 0805 hr on 22 June.

The drill bit/mud motor and casing shoe encountered some resistance passing through the reentry cone hanger at 4686 mbrf. After lowering the drill string to 4842 mbrf (156 mbsf), we began washing and drilling in the ACORK, reaching 5269 mbrf (583 mbsf) at 2400 hr on 22 June. Twenty-barrel mud (sepiolite) sweeps were circulated at 295, 430, 497, 555, and 574 mbsf. The drilling parameters were a WOB of 0–20 klb with a circulation rate of 100–120 spm and a pressure of 1450–2200 psi.

On 23 June, we continued drilling in the ACORK from 5269 mbrf (583 mbsf) to 5452 mbrf (766 mbsf). The VIT camera had to be retrieved at 2030 hr. Twenty-barrel mud (sepiolite) sweeps were circulated at 660, 670, 747, and 764 mbsf. The penetration rate for 22 June was 7.6 m/hr. The drilling parameters were a WOB of 10–20 klb and a circulation of 120 spm with a pressure of 2200 psi.

From 24 to 26 June, the ACORK was advanced from 5452 mbrf (766 mbsf) to 5611 mbrf (925 mbsf) with a penetration rate of only ~2 m/hr. Twenty-barrel mud (sepiolite) sweeps were circulated every 10–20 m. The drilling parameters were a WOB ranging up to ~180 klb and a circulation of 70–120 spm with a pressure up to 2200 psi.

As the penetration rate was so low and any increase in WOB would cause concern for the safety of the casing and drill string, we decided to deploy the VIT camera at 1400 hr on 26 June and perform some WOB tests. Once the camera was on the bottom at 1615 hr, we noted that the drill string, ACORK head, casing, and reentry cone all appeared to be intact. We then attempted to increase the WOB to the maximum deemed safe and saw no significant increase in penetration or bending of the drill string or ACORK. We also tried to raise the drill string to see if recovering the entire assembly was an option; it could not be raised. During these tests, circulation was observed coming out of the reentry cone, which indicated that, although a flow path existed, the casing was being held by skin friction. The VIT was retrieved at 1950 hr on 26 June.

Because it was clear that we could not pull out of the hole, we continued to try to advance the ACORK in the hope that we could advance enough to either latch it in the reentry cone or get it low enough so that it could remain upright on its own. On 27 June, we advanced from 5611 mbrf (925 mbsf) to 5620 mbrf (934 mbsf) or ~0.4 m/hr. Two 100-bbl mud sweeps were pumped at 927 and 929 mbsf, as were two 20-bbl sweeps at 931 and 933 mbsf. The drilling parameters were a WOB of 180 klb and a pump rate of 70–120 spm with a pressure of 1000–2200 psi

From 0000 to 0430 hr on 28 June, we continued to work the drill string, but it would not advance below 5620 mbrf (934 mbsf). In a last-ditch effort, we pumped 500-bbl of heavy (11.3 lb/gal) mud into the hole at 0430 hr and continued to work the pipe until 0730 hr with no success. At this point, we decided to inflate the packers. The circulation sub was shifted open with the wireline-shifting tool. The go-devil was dropped; once it landed, a slow pump rate was used to inflate the two packers and close the spool valves. This was completed at 0900 hr.

We then released the casing running tool at 1100 hr and raised the drill string from 5620 mbrf (934 mbsf) to 5574 mbrf (888 mbsf). Because it seemed unlikely that conditions would allow a bridge plug to be installed, the science party requested heavy mud be placed in the hole, so 100 bbl (10.5 lb/gal) of mud was pumped into the hole. Tripping operations resumed at 1215 hr, with the top drive being racked back and the 20-ft knobby laid out.

The VIT camera was deployed and run in the hole from 1245 to 1315 hr on 28 June to observe the withdrawal of the drilling BHA from the ACORK head. At 1533 hr the drilling BHA came out of the reentry cone, and the ACORK head slowly bent toward the seafloor and went out of sight. The drill string and VIT camera were lowered for a closer inspection, and the ACORK casing was observed to be bent over the edge of the reentry cone. With the camera, we followed the casing away from the reentry cone. The compass on the camera indicated the casing fell to the north. No damage was seen to the ACORK umbilical, casing, or head. The ACORK head was lying flat on the seafloor, and ~60% of the ACORK head appeared above the seafloor. Most importantly, the underwater mateable connector was sticking straight up so that a submarine or ROV can still connect to download the data.

We started to recover the VIT camera at 1645 hr on 28 June. The drill string was pulled to 4491 mbrf, and knobbies were installed so that the drill string could be hung off during the DP transit back to Hole 1173B. The seafloor positioning beacon was released at 1815 hr. Once the VIT camera was back on board at 1845 hr, the hydrophones were raised, and the ship began the DP transit back to Hole 1173B.

Next Section | Table of Contents